# Geometric foundations of thermodynamics in the quantum regime

### v1

@@ XI.2 Curvature
$\omega=\sum_{\alpha=0}^{n}\sum_{k=1}^{n}\left(dv_{\alpha}-\Gamma^{k}_{\alpha}d\lambda_{k}\right)\otimes\dfrac{\partial}{\partial v_{\alpha}}.$ (77)
This is the pullback of the Maurer-Cartan form minus the gauge potential $\Gamma^{k}_{\alpha}$. The coefficients $\Gamma^{k}_{\alpha}(\boldsymbol{\lambda})$ encode the geometric coupling between changes in intensive parameters $\boldsymbol{\lambda}$ and shifts in thermodynamic labels $(S,\boldsymbol{a})$ during parallel transport. The connection coefficients $\Gamma^{k}_{\alpha}$ are determined by the orthogonality condition $g_{\mathcal{M}}(e_{k},\partial_{v_{\beta}})=0$. Explicitly:
  • For $\beta=0$, i.e. $\partial_{v_{0}}=\partial_{S}$:
    $g_{\mathcal{M}}\left(e_{k},\dfrac{\partial}{\partial S}\right)=g_{\mathcal{M}}\left(\dfrac{\partial}{\partial\lambda_{k}}-\sum_{\alpha=0}^{n}\Gamma^{k}_{\alpha}\dfrac{\partial}{\partial v_{\alpha}},\dfrac{\partial}{\partial S}\right)=g_{S\lambda_{k}}-\Gamma^{k}_{0}g_{SS}=h_{k}-\Gamma^{k}_{0}g_{S}=0.$ (78)
    Thus, $\Gamma^{k}_{0}=\frac{h_{k}}{g_{S}}$.
  • For $\beta=i\in\{1,\dots,n\}$, i.e., $\partial_{v_{i}}=\partial_{a_{i}}$:
    $\begin{split}g_{\mathcal{M}}\left(e_{k},\dfrac{\partial}{\partial a_{i}}\right)&=g_{\mathcal{M}}\left(\dfrac{\partial}{\partial\lambda_{k}}-\sum_{\alpha=0}^{n}\Gamma^{k}_{\alpha}\dfrac{\partial}{\partial v_{\alpha}},\dfrac{\partial}{\partial a_{i}}\right)=-\Gamma^{k}_{i}g_{a_{i}a_{i}}=-\Gamma^{k}_{i}g_{a_{i}}=0.\end{split}$ (79)
@@
$\Gamma^{k}_{\alpha}(\boldsymbol{\lambda})=\begin{cases}\displaystyle h_{k}(\boldsymbol{\lambda})/g_{S}(\boldsymbol{\lambda})&\alpha=0\\ 0&\alpha=1,\dots,n\end{cases}$ (80)
@@
The connection acts exclusively in the entropy direction, reflecting that changes in intensive parameters $\boldsymbol{\lambda}$ induce adjustments in thermodynamic entropy but not in target expectation values, consistent with the structure of Gibbs states.
@@
$e_{k}=\dfrac{\partial}{\partial\lambda_{k}}-\sum_{\alpha=0}^{n}\Gamma^{k}_{\alpha}\dfrac{\partial}{\partial v_{\alpha}},$ (81)
@@
$R(e_{k},e_{l})=\sum_{\alpha=0}^{n}\left(\dfrac{\partial}{\partial\lambda_{k}}\Gamma^{l}_{\alpha}-\dfrac{\partial}{\partial\lambda_{l}}\Gamma^{k}_{\alpha}\right)\dfrac{\partial}{\partial v_{\alpha}}.$ (82)
@@
$\omega=\sum_{\alpha=0}^{n}\sum_{m=1}^{n}\left(dv_{\alpha}-\Gamma^{m}_{\alpha}d\lambda_{m}\right)\otimes\dfrac{\partial}{\partial v_{\alpha}}.$ (83)
@@
$\begin{split}d\omega&=\sum_{\alpha=0}^{n}\sum_{k=1}^{n}\sum_{l=1}^{n}\left(\dfrac{\partial}{\partial\lambda_{k}}\Gamma^{k}_{\alpha}\right)d\lambda_{l}\wedge d\lambda_{k}\otimes\dfrac{\partial}{\partial v_{\alpha}}\\ &=\sum_{\alpha=0}^{n}\sum_{k=1}^{n}\sum_{l=1}^{n}\frac{1}{2}\left(\dfrac{\partial}{\partial\lambda_{k}}\Gamma^{l}_{\alpha}-\dfrac{\partial}{\partial\lambda_{l}}\Gamma^{k}_{\alpha}\right)d\lambda_{k}\wedge d\lambda_{l}\otimes\dfrac{\partial}{\partial v_{\alpha}}.\end{split}$ (84)
@@
$R(e_{k},e_{l})=d\omega(e_{k},e_{l})=\sum_{\alpha=0}^{n}\left(\dfrac{\partial}{\partial\lambda_{k}}\Gamma^{l}_{\alpha}-\dfrac{\partial}{\partial\lambda_{l}}\Gamma^{k}_{\alpha}\right)\dfrac{\partial}{\partial v_{\alpha}},$ (85)
which is Eq. (82). Note that this expression is similar to the general one given by Eq. (62). The Lie bracket vanishes because the group is abelian.∎
Substituting the explicit connection coefficients from Eq. (80), we have that only the $\alpha=0$ term contributes to the curvature:
$R(e_{k},e_{l})$ $=\left[\dfrac{\partial}{\partial\lambda_{k}}\left(\frac{h_{l}}{g_{S}}\right)-\dfrac{\partial}{\partial\lambda_{l}}\left(\frac{h_{k}}{g_{S}}\right)\right]\dfrac{\partial}{\partial S}.$
@@
Curvature acts exclusively in the entropy direction, $\partial_{S}$, measuring the non-integrability of the entropy compensation induced by changes in the intensive parameters $\boldsymbol{\lambda}$.
@@
$\dfrac{\partial}{\partial\lambda_{k}}\left(\frac{h_{l}}{g_{S}}\right)=\dfrac{\partial}{\partial\lambda_{l}}\left(\frac{h_{k}}{g_{S}}\right),$ (86)
for all $k,l$, i.e. the $1$-form
$\psi_{i}=\dfrac{h_{i}}{g_{S}}d\lambda_{i},$ (87)
@@
a vertical displacement in the fiber. Non-zero holonomy implies that parallel transport along a cycle in the space of Gibbs states shifts the thermodynamic coordinates $(S,\boldsymbol{a})$, inducing geometric irreversibility $\Sigma_{\gamma}>0$ unless $R=0$.
In this gauge framework, thermodynamic processes are sections of associated vector bundles, and quasistatic transformations correspond to paths with minimal holonomic deviation. Irreversibility in cyclic processes emerges as a curvature effect, unifying thermodynamic dissipation with geometric phases in gauge theories. This mirrors the Aharonov-Bohm phase in electromagnetism or Berry phase in quantum mechanics, where holonomy in a principal bundle over parameter space leads to observable shifts despite local flatness. Here, non-zero $R$ generates non-integrable shifts in $(S,\boldsymbol{a})$, requiring dissipative corrections to close the cycle, manifesting as $\Sigma_{\gamma}>0$—a gauge-induced irreversibility intrinsic to the thermodynamic bundle geometry.
If the principal connection on the quantum thermodynamic bundle is flat, then all cyclic processes in the base manifold $\mathcal{B}$, corresponding to closed loops in intensive parameters $\boldsymbol{\lambda}$, induce zero holonomy. This implies that parallel transport along such cycles returns the system to its initial thermodynamic configuration without any geometric shift in labels $(S,\boldsymbol{a})$, implying no additional entropy production from bundle geometry. In practical terms, flat connections characterize systems where thermodynamic cycles have no gauge-like losses, provided the parameter space $\mathcal{B}$ possesses trivial topology.
@@
The quantum second law emerges from the non-integrability of the contact structure and the geometry of non-equilibrium paths. For a finite-speed process along a path $\gamma(t)$ in $\mathcal{E}$, the entropy production rate is $\varsigma(t)=\kappa g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))\geq 0$, with total $\Sigma=\kappa\int_{0}^{T}g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))dt\geq 0$. Irreversibility is minimal along geodesics and vanishes in the quasistatic limit. Off equilibrium, in fibers $F_{\sigma}$, relaxation paths incur positive entropy production $\Sigma>0$ due to deviations from $\mathcal{E}$. Curvature-induced holonomy in cyclic processes generates additional geometric entropy production.
Remark 21 (Classical thermodynamics as a trivial limit).

### v2

@@ XI.2 Curvature on the principal bundle
$\omega=\sum_{\alpha=0}^{n}\sum_{k=1}^{n}\left(dv_{\alpha}-\Gamma^{k}_{\alpha}d\bar{\lambda}_{k}\right)\otimes\dfrac{\partial}{\partial v_{\alpha}}.$ (87)
This is the pullback of the Maurer-Cartan form minus the gauge potential $\Gamma^{k}_{\alpha}$. The coefficients $\Gamma^{k}_{\alpha}(\bar{\boldsymbol{\lambda}})$ encode the geometric coupling between changes in intensive parameters $\bar{\boldsymbol{\lambda}}$ and shifts in thermodynamic labels $(\bar{S},\bar{\boldsymbol{a}})$ during parallel transport. The connection coefficients $\Gamma^{k}_{\alpha}$ are determined by the orthogonality condition $g_{\mathcal{M}}(e_{k},\partial_{v_{\beta}})=0$. Explicitly:
  • For $\beta=0$, i.e., $\partial_{v_{0}}=\partial_{\bar{S}}$:
    $g_{\mathcal{M}}\left(e_{k},\dfrac{\partial}{\partial\bar{S}}\right)=g_{\mathcal{M}}\left(\dfrac{\partial}{\partial\bar{\lambda}_{k}}-\sum_{\alpha=0}^{n}\Gamma^{k}_{\alpha}\dfrac{\partial}{\partial v_{\alpha}},\dfrac{\partial}{\partial\bar{S}}\right)=g_{\bar{S}\bar{\lambda}_{k}}-\Gamma^{k}_{0}g_{\bar{S}\bar{S}}=h_{k}-\Gamma^{k}_{0}g_{\bar{S}}=0.$ (88)
    Thus, $\Gamma^{k}_{0}=h_{k}/g_{\bar{S}}$.
  • For $\beta=i\in\{1,\dots,n\}$, i.e., $\partial_{v_{i}}=\partial_{\bar{a}_{i}}$:
    $\begin{split}g_{\mathcal{M}}\left(e_{k},\dfrac{\partial}{\partial\bar{a}_{i}}\right)&=g_{\mathcal{M}}\left(\dfrac{\partial}{\partial\bar{\lambda}_{k}}-\sum_{\alpha=0}^{n}\Gamma^{k}_{\alpha}\dfrac{\partial}{\partial v_{\alpha}},\dfrac{\partial}{\partial\bar{a_{i}}}\right)=-\Gamma^{k}_{i}g_{\bar{a}_{i}\bar{a}_{i}}=-\Gamma^{k}_{i}g_{\bar{a}_{i}}=0.\end{split}$ (89)
@@
$\Gamma^{k}_{\alpha}(\boldsymbol{\bar{\lambda}})=\begin{cases}\displaystyle h_{k}(\bar{\boldsymbol{\lambda}})/g_{\bar{S}}(\bar{\boldsymbol{\lambda}})&\text{if }\alpha=0,\\ 0&\text{if }\alpha=1,\dots,n.\end{cases}$ (90)
@@
Note that the connection acts exclusively in the entropic direction, reflecting that changes in intensive parameters $\bar{\boldsymbol{\lambda}}$ induce adjustments in the value of $\bar{S}$ but not in the value of the coordinates $\bar{a}_{i}$, consistent with the structure of the quantum thermodynamic fiber bundle.
@@
$e_{k}=\dfrac{\partial}{\partial\bar{\lambda}_{k}}-\sum_{\alpha=0}^{n}\Gamma^{k}_{\alpha}\dfrac{\partial}{\partial v_{\alpha}},$ (91)
@@
$R(e_{k},e_{l})=\sum_{\alpha=0}^{n}\left(\dfrac{\partial}{\partial\bar{\lambda}_{k}}\Gamma^{l}_{\alpha}-\dfrac{\partial}{\partial\bar{\lambda}_{l}}\Gamma^{k}_{\alpha}\right)\dfrac{\partial}{\partial v_{\alpha}}.$ (92)
@@
$\omega=\sum_{\alpha=0}^{n}\sum_{m=1}^{n}\left(dv_{\alpha}-\Gamma^{m}_{\alpha}d\bar{\lambda}_{m}\right)\otimes\dfrac{\partial}{\partial v_{\alpha}}.$ (93)
@@
$\begin{split}d\omega&=\sum_{\alpha=0}^{n}\sum_{k=1}^{n}\sum_{l=1}^{n}\left(\dfrac{\partial}{\partial\bar{\lambda_{k}}}\Gamma^{k}_{\alpha}\right)d\bar{\lambda}_{l}\wedge d\bar{\lambda}_{k}\otimes\dfrac{\partial}{\partial v_{\alpha}}\\ &=\sum_{\alpha=0}^{n}\sum_{k=1}^{n}\sum_{l=1}^{n}\frac{1}{2}\left(\dfrac{\partial}{\partial\bar{\lambda}_{k}}\Gamma^{l}_{\alpha}-\dfrac{\partial}{\partial\bar{\lambda}_{l}}\Gamma^{k}_{\alpha}\right)d\bar{\lambda}_{k}\wedge d\bar{\lambda}_{l}\otimes\dfrac{\partial}{\partial v_{\alpha}}.\end{split}$ (94)
@@
$R(e_{k},e_{l})=d\omega(e_{k},e_{l})=\sum_{\alpha=0}^{n}\left(\dfrac{\partial}{\partial\bar{\lambda}_{k}}\Gamma^{l}_{\alpha}-\dfrac{\partial}{\partial\bar{\lambda}_{l}}\Gamma^{k}_{\alpha}\right)\dfrac{\partial}{\partial v_{\alpha}},$ (95)
which is Eq. (92). Note that this expression is similar to the general one given by Eq. (72). The Lie bracket vanishes since the group is abelian. ∎
Substituting the explicit connection coefficients from Eq. (90), we have that only the $\alpha=0$ term contributes to the curvature
$R(e_{k},e_{l})$ $=\left[\dfrac{\partial}{\partial\bar{\lambda}_{k}}\left(\frac{h_{l}}{g_{\bar{S}}}\right)-\dfrac{\partial}{\partial\bar{\lambda}_{l}}\left(\frac{h_{k}}{g_{\bar{S}}}\right)\right]\dfrac{\partial}{\partial\bar{S}}.$ (96)
Thus, $R$ is a $\mathfrak{g}$-valued 2-form acting solely in the entropic direction, reflecting that non-commutativities in intensive parameter variations induce non-integrable shifts in thermodynamic entropy.
Remark 16.
The curvature $R$ of the principal connection is a $\mathfrak{g}$-valued $2$-form on $\mathcal{M}$, where $\mathfrak{g}=\mathbb{R}^{n+1}$ is the Lie algebra of the structure group $G=(\mathbb{R}^{n+1},+)$. In the local trivialization, it takes the form
$R=\frac{1}{2}\sum_{k,l=1}^{n}\left[\frac{\partial}{\partial\bar{\lambda}_{k}}\left(\frac{h_{l}}{g_{\bar{S}}}\right)-\frac{\partial}{\partial\bar{\lambda}_{l}}\left(\frac{h_{k}}{g_{\bar{S}}}\right)\right]d\bar{\lambda}_{k}\wedge d\bar{\lambda}_{l}\otimes\frac{\partial}{\partial\bar{S}}.$ (97)
Since the structure group is abelian, $R=d\omega$. The tensor product with $\partial_{\bar{S}}$ reflects that the curvature takes values exclusively in the entropic component of $\mathfrak{g}$; all components in the $\partial_{\bar{a}_{i}}$ directions vanish due to the absence of cross terms coupling $d\bar{a}_{i}$ and $d\bar{\lambda}_{j}$ in the metric $g_{\mathcal{M}}$. The expression given by Eq. (97) is the standard coordinate expression of a Lie-algebra-valued $2$-form (see e.g. [32, 11]).
@@
Curvature acts exclusively in the entropic direction, $\partial_{\bar{S}}$, measuring the non-integrability of the entropy compensation induced by changes in the intensive parameters $\bar{\boldsymbol{\lambda}}$.
@@
$\dfrac{\partial}{\partial\bar{\lambda}_{k}}\left(\frac{h_{l}}{g_{\bar{S}}}\right)=\dfrac{\partial}{\partial\bar{\lambda}_{l}}\left(\frac{h_{k}}{g_{\bar{S}}}\right),$ (98)
for all $k,l$, i.e., the $1$-form
$\psi_{i}=\dfrac{h_{i}}{g_{\bar{S}}}d\bar{\lambda}_{i},$ (99)
@@
a vertical displacement in the fiber. Non-zero holonomy implies that parallel transport along a cycle in the space of Gibbs states shifts the thermodynamic coordinates $(\bar{S},\bar{\boldsymbol{a}})$, inducing geometric irreversibility $\Sigma\neq 0$ unless $R=0$.
In this gauge framework, quasistatic transformations correspond to paths with minimal holonomic deviation. Irreversibility in cyclic processes emerges as a curvature effect, unifying thermodynamic dissipation with geometric phases in gauge theories. This mirrors the Aharonov-Bohm phase in electromagnetism or Berry’s phase in quantum mechanics, where holonomy in a principal bundle over parameter space leads to observable shifts despite local flatness. Here, non-zero $R$ generates non-integrable shifts in $(\bar{S},\bar{\boldsymbol{a}})$, requiring dissipative corrections to close the cycle, manifesting as $\Sigma\neq 0$—a gauge-induced irreversibility intrinsic to the thermodynamic bundle geometry.
If the principal connection on the quantum thermodynamic bundle is flat, then all cyclic processes in the base manifold $\mathcal{B}$, corresponding to closed loops in intensive parameters $\bar{\boldsymbol{\lambda}}$, induce zero holonomy. This implies that parallel transport along such cycles returns the system to its initial thermodynamic configuration without any geometric shift in labels $(\bar{S},\bar{\boldsymbol{a}})$, implying no additional entropy production from bundle geometry. In practical terms, flat connections characterize systems whose thermodynamic cycles have no gauge-like losses, provided the parameter space $\mathcal{B}$ possesses trivial topology.
The principal fiber bundle structure introduces gauge symmetry into quantum thermodynamics. The additive group acts by translating the coordinates $(\bar{S},\bar{\boldsymbol{a}})$ while preserving the quantum state and intensive parameters, representing a redundancy in thermodynamic descriptions akin to gauge freedoms in field theories where different potentials give the same physical fields. Equilibrium serves as a gauge-fixing condition, ensuring a unique canonical representation per state due to the maximum entropy principle. The connection 1-form $\omega$ acts as a gauge field, coupling changes in parameters to label shifts, with (in this case) its abelian curvature $R=d\omega$ (or $d\omega+\omega\wedge\omega$ in the non-abelian case) focused on the entropy direction, measuring non-commutativity in parameter evolutions. As mentioned in the previous section, the direct consequences include holonomy-imposed bounds on quantum engine efficiencies, where cyclic protocols accumulate unremovable dissipation. This opens the avenue for optimization of control protocols exploiting gauge choices, and conceptual parallels to thermodynamic analogs of the Aharonov-Bohm effect, potentially inspiring experimental probes of geometric phases in thermal machines.
@@
The quantum second law emerges from the non-integrability of the contact structure and the geometry of non-equilibrium paths. For a finite-speed process along a path $\gamma(t)$ in $\mathcal{E}$, the entropy production rate is $\varsigma(t)=\kappa g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))\geq 0$, with total $\Sigma=\kappa\int_{0}^{T}g_{\mathrm{BW}}(\gamma^{\prime}(t),\gamma^{\prime}(t))dt\geq 0$. Irreversibility is minimal along geodesics and vanishes in the quasistatic limit. Off equilibrium, in fibers $F_{\sigma}$ and assuming that the full bundle metric $g_{\mathcal{M}}$ is Riemannian, relaxation paths incur positive entropy production $\Sigma>0$ due to deviations from $\mathcal{E}$. Curvature-induced holonomy in cyclic processes generates additional geometric entropy production.
XI.3 Gauge and physical interpretation of curvature
The curvature of the principal connection on the quantum thermodynamic fiber bundle is the standard curvature $2$-form of mathematical gauge theory, as defined in standard references such as [22, 32, 11]. This object admits several complementary physical interpretations that clarify its thermodynamic significance.
First, in the Ehresmann picture, the curvature measures the failure of horizontal integrability. Specifically, $R(X,Y)=\operatorname{\mathrm{proj}}_{\mathscr{V}}([X,Y])$ for horizontal vector fields $X,Y\in\mathfrak{X}(\mathscr{H})$ quantifies how much parallel transport around an infinitesimal closed loop fails to return to the initial point in the fiber due to the non-integrability of the horizontal distribution $\mathscr{H}$. In local coordinates, $R(e_{k},e_{l})$ is the vertical projection of the Lie bracket of the horizontal lifts $e_{k},e_{l}$. Non-zero curvature then implies that thermodynamic labels $(\bar{S},\bar{\boldsymbol{a}})$ do not return to their initial values after a closed cycle in intensive-parameter space $\bar{\boldsymbol{\lambda}}$. This is the primary geometric mechanism underlying holonomy-induced irreversibility in the framework.
In the language of gauge field theory, the curvature $R\equiv F=d\omega+\omega\wedge\omega$ is equivalent to the strength of the local gauge field. In a local trivialization, $R$ acts as the force field felt by the thermodynamic coordinates when transported along the base manifold. In the abelian case, $R$ takes values exclusively in the entropy direction $\partial_{\bar{S}}$, inducing a net shift $\Delta\bar{S}$ during cyclic variations of $\bar{\boldsymbol{\lambda}}$—analogous to how the electromagnetic field strength $F_{\mu\nu}$ generates forces via the covariant derivative $D=d+A$ whose square is $D^{2}=F$.
From a topological viewpoint, $R$ is the source of holonomy. Parallel transport around a closed loop $\gamma$ in the base generates the holonomy $\mathrm{Hol}(\gamma,p_{0})=\mathcal{P}\exp(-\oint_{\gamma}\omega)$, which, in the abelian case and by Stokes’ theorem, reduces to $\exp(-\int_{\mathscr{S}}R)$ for any surface $\mathscr{S}$ bounding $\gamma$. This holonomic shift in the thermodynamic coordinates, typically entropy, requires dissipative corrections to restore equilibrium, providing a purely geometric origin for irreversibility in cyclic processes. This interpretation is the direct thermodynamic counterpart of the Berry phase in quantum mechanics and the Aharonov–Bohm effect in gauge theory. By the Ambrose–Singer theorem, the curvature generates the holonomy group: the Lie algebra of the holonomy group is spanned by the values of $R$ evaluated on horizontal vectors. Non-zero curvature thus produces a non-trivial holonomy group acting on the fiber, manifesting as topological irreversibility in cyclic thermodynamic protocols.
Finally, $R$ is simply a Lie algebra-valued $2$-form $R\in\Lambda^{2}(\mathcal{M};\mathfrak{g})$ with $\mathfrak{g}=\mathbb{R}^{n+1}$, encoding the (abelian) non-commutativity of infinitesimal gauge transformations. In potential non-abelian extensions (e.g., systems with internal symmetries or constrained ensembles), the $\omega\wedge\omega$ term would appear, leading to non-commuting holonomies consistent with Wilczek–Zee phases, see Section XIII.2.
XI.4 Toward topological quantum thermodynamics: invariants and topological considerations in the quantum thermodynamic bundle
Following the exposition in Ref. [22], in the principal bundle $(\mathcal{M},\mathcal{B},\Xi,\mathbb{R}^{n+1})$, quantities unchanged under bundle automorphisms, i.e., arbitrary re-gaugings of the thermodynamic labels $(\bar{S},\bar{\boldsymbol{a}})$ while keeping the physical state $\sigma\in\mathcal{B}$ fixed, are gauge invariants. These invariants encode the intrinsic geometry and topology of the connection $\omega$ and its curvature $R$, furnishing universal, gauge-independent lower bounds on irreversible entropy production in cyclic quantum thermodynamic processes.
Differences of connections $\omega_{1}-\omega_{2}$ transform as 1-forms on $\mathcal{B}$, forming an affine space of connections. The curvature descends to a well-defined closed 2-form $R_{\mathcal{B}}\in\Lambda^{2}(\mathcal{B};\mathbb{R}^{n+1})$ on the base, quantifying label shifts independent of the gauging. Bianchi’s identity $dR+[\omega,R]=0$, where $dR=0$ for the abelian case, ensures closure of $R_{\mathcal{B}}$. That is, since the structure group is abelian, Bianchi’s identity reduces to $dR_{\mathcal{B}}=0$, implying that the holonomy around any closed loop $\gamma\subset\mathcal{B}$ depends only on its homotopy class $\gamma]\in\pi_{1}(\mathcal{B})$. Consequently, topologically equivalent cycles induce identical geometric entropy shifts.
A natural gauge-invariant diagnostic for a loop $\gamma:S^{1}\to\mathcal{B}$ is the Wilson loop on an associated vector bundle $E=\mathcal{M}\times_{\rho}V$
$W^{E}_{\gamma}(\omega)=\mathrm{tr}\left[\mathcal{P}\exp\left(-\oint_{\gamma}\rho_{*}\omega^{s}\right)\right],$ (101)
where $\omega^{s}=s^{*}\omega$ is the pull-back of the connection in a local trivialization over an open set containing $\gamma$ with global gauge $s:U\to\mathcal{M}$; and $V$ is the internal vector space on which the connection acts through the representation $\rho:\mathbb{R}^{n+1}\to\mathrm{GL}(V)$, the trivial representation on the fiber of $E$.
In the abelian case, path-ordering reduces to the ordinary exponential and the trace is unnecessary for line bundles, corresponding to scalar holonomy shifts in labels, so
$W^{E}_{\gamma}(\omega)=\exp\left(-\oint_{\gamma}\omega^{s}\right).$ (102)
When projected onto the entropy component of the fiber, this gives a real number $\Delta\bar{S}_{\mathrm{geo}}$, measuring the net shift in thermodynamic labels after parallel transport. This geometric contribution adds to the total entropy production $\Sigma$ in cyclic processes, providing a strict lower bound on dissipation for any cyclic protocol.
Topologically, with an Ad-invariant scalar product $\langle\cdot,\cdot\rangle$ on $\mathfrak{g}=\mathbb{R}^{n+1}$, e.g., the Euclidean scalar product, the Chern–Simons 3-form on $\mathcal{M}$ is defined as
$\alpha(\omega)=\langle\omega,R\rangle\in\Lambda^{3}(\mathcal{M}),$ (103)
satisfying $d\alpha=\langle R,R\rangle$. Here, $\langle R,R\rangle$ acts as a local geometric measure of dissipation induced by curvature. Under automorphisms, $\alpha$ transforms by an exact term plus a closed 3-form from the structure group, so the integral
$\mathcal{S}_{\mathrm{CS}}(\omega)=\int_{\mathscr{M}}\alpha(\omega)\pmod{\mathbb{Z}}$, (104)
over closed oriented 3-submanifolds $\mathscr{M}\subset\mathcal{M}$, e.g., a region swept by a continuous family of loops in parameter space, is gauge-invariant and metric-independent. This number is known as Chern–Simons action. A non-vanishing value highlights the existence of topologically protected fractional contributions to entropy production: quasistatic cycles enclosing non-contractible 3-volumes cannot be made dissipation-free, no matter how slowly they are driven. Note that this Chern–Simons invariant is purely topological, and does not depend on the metric chosen.
Flat connections give $\mathcal{S}_{\mathrm{CS}}=0$, permitting reversible cycles. Differences $\omega-\omega_{0}$ transform adjointly, and $\mathcal{S}_{\mathrm{CS}}$ classifies topological sectors—universal bounds on dissipation in quantum engines over non-simply connected parameter spaces, e.g., periodic $\boldsymbol{\lambda}$ resulting in toroidal $\mathcal{B}$.
In natural extensions where the abelian group is compactified, e.g., $\mathbb{R}\to S^{1}$ arising from periodic intensive parameters or modular identifications, additional topological invariants emerge from characteristic classes. For a principal $S^{1}$-bundle $\mathcal{M}\to\mathcal{B}$ equipped with the connection $\omega$, the curvature defines the first Chern class
$c_{1}(\mathcal{M})=-\frac{1}{2\pi i}[R_{\mathcal{B}}]\in H^{2}_{\mathrm{dR}}(\mathcal{B};\mathbb{R}),$ (105)
a real de Rham cohomology class independent of the choice of connection. A non-trivial $c_{1}$ precludes the existence of a global section, that is, the existence of a unique global thermodynamic labeling. It also forces irreducible holonomy around non-contractible loops, even if the curvature vanishes locally.
Example 4.
In particular, consider the canonical Hopf bundle $S^{1}\hookrightarrow S^{3}\xrightarrow{\pi}S^{2}\cong\mathbb{CP}^{1}$ [23]. The total space is defined as the unit sphere in $\mathbb{C}^{2}$, $S^{3}=\{(z_{1},z_{2})\in\mathbb{C}^{2}||z_{1}|^{2}+|z_{2}|^{2}=1\}$. The projection map $\pi:S^{3}\to\mathbb{CP}^{1}$ sends each point to the complex line it spans. The fiber over each point of the base $S^{2}$ is a circle $S^{1}$ consisting of all global phases $e^{i\theta}(z_{1},z_{2})$ consistent with a given projective point in $\mathbb{CP}^{1}$. By taking the curvature 2-form $R$ corresponding to a Dirac monopole field of unit strength, we can generate the fundamental class of $H^{2}(S^{2};\mathbb{Z})\cong\mathbb{Z}$. Therefore, the first Chern class is represented by
$c_{1}(\mathrm{Hopf})=\left[\frac{R}{2\pi}\right],$ (106)
and integrating it over the base space gives
$\int_{S^{2}}c_{1}(\mathrm{Hopf})=1.$ (107)
Note that $\pi_{1}(S^{2})=0$: every closed loop in $S^{2}$ is contractible. From the long exact sequence of homotopy groups of the fibration, the connecting homomorphism yields an isomorphism $\pi_{2}(S^{2})\cong\pi_{1}(S^{1})\cong\mathbb{Z}$. This shows that the non-triviality of the bundle is inherently tied to the second homotopy group wrapping around the fiber, which is equivalently measured by the first Chern class. In our thermodynamic interpretation, this integer $1$ is the minimal topological charge: any cyclic driving protocol whose control manifold contains a 2-cycle homologous to the generator of $H_{2}(\mathcal{B})$ must induce at least one unit of topological entropy shift $\Delta\bar{S}_{\mathrm{geo}}=1$.
For non-abelian compactifications, e.g., SU(2) bundles for symmetry-protected degeneracies in spin systems, higher Chern classes $c_{k}(P)\in H^{2k}_{\mathrm{dR}}(\mathcal{B})$ impose further constraints on geometric dissipation through integrated curvature.
Therefore, all these invariants unify topological dissipation. Wilson loops bound local cycle production, while Chern–Simons integrals and Chern classes set universal lower limits on entropy production in topologically non-trivial control protocols. Geometric irreversibility is thereby elevated to a gauge- and topology-protected phenomenon.
XI.5 Classical limit